OpenStudy (anonymous):

Solve the equation -2(5-3n)+8=4+5n A. -6/11 B. 1/6 C. 4/9 D.6

OpenStudy (anonymous):

I got 6

OpenStudy (jessicap):

-2(5-3n)+8=4+5n 1. Distribute the 2 -10+6n+8=4+5n 2. Simplify 6n-2=4+5n 3. Combine like terms 6n-2+2=4+5n+2 4. 2 cancels on the left side, add to 4 on the right 6n=6+5n 5. Subtract 5n from both sides n=6 The answer is D (idk where -22 came from :)